Find the amount of water at 20°c required to just melt 5g of ice at -20°c

Grade:11

2 Answers

Arun
6 years ago
Dear Sagar
 
m*4.18 * 20 = 5 *80 + 5 * 0.5 * 20
4.18 m = 20 + 2.5
m = 22.5/4.18 = 5.38
 
Regards
Arun (askIITians forum expert)
Prajakta tilekar
6 years ago
At -20c ice absorb 50 cal and at 20c water released 20x if we consider mass as x. so at 0c ice should absorb 20x-50 cal. So 20x-50 =Mmelt×80 so mmelt is 20x-50/80 and we require total mass of ice should be melt so 20x-50/80=5 we get x=22.5